# From shape to fate: making bacterial swarming expansion predictable

### v1

@@ 2.2 Recovering colony fronts for accurate morphological forecasting
Figure 3: TexPol–Net improves colony-front segmentation by coupling texture-sensitive boundary encoding with a geometry-aligned context prior. a, Texture-Polar Network (TexPol–Net) within a prototype-based instance segmentation pipeline. Texture-Edge Attention (TEA) is embedded in the backbone to preserve high-frequency boundary cues, while Polar–Context Attention (PCA) is interleaved in the PANet-style bidirectional neck to maintain polar consistency during multi-scale fusion. Dense heads at P3P3P5P5 predict class scores, boxes and mask coefficients; a lightweight Protonet generates k=32k{=}32 shared prototypes that are linearly combined into instance masks and post-processed by cropping and thresholding. b, Qualitative comparison on two representative swarming regimes. In the dense–finger phase (top), YOLOv11 [19] captures the coarse outline but smooths the corrugated front, shortening narrow fingers and reducing anisotropy. SAM [20] and SAM2 [34] favor the colony core, producing a boundary that approaches a near-circular shape. TexPol–Net preserves finger heterogeneity and radial bias, with YOLOv12 [41] as the closest competing backbone. In the near–concentric ring regime (bottom), all methods recover the circular structure, while TexPol–Net more precisely localizes the advancing boundary. c, Average precision as a function of IoU matching threshold, highlighting boundary-stringent performance at high IoU. d, Image-wise overlap statistics for IoU and Dice, summarizing accuracy and variability across images.
Accurate recovery of the advancing front is a structural prerequisite for morphological forecasting rather than a superficial preprocessing step. Errors introduced at the segmentation stage propagate directly into the temporal model, as smoothing finger-like protrusions, displacing the outer rim, or introducing spurious fragments alters the effective front trajectories available for learning. By preserving fine-scale protrusions while maintaining the global organization of the rim, TexPol–Net yields a geometrically consistent sequence of morphological masks in which successive fronts are linked through physically plausible deformations. This representation defines an appropriate input substrate for curvature-driven expansion and underpins the forecasting analyses presented in the following sections.
2.3 Forecasting the expanding front of swarming colonies
We conduct a systematic comparison between Morpher and representative state–of–the–art video prediction models, including MAU [6], MIM [46], PredRNN [44], PredRNNv2 [45], the original SimVP implementation which employs the TAU temporal unit [15, 40], and the improved SimVPv2 variant, in which TAU is replaced by the gSTA module [39]. These models have demonstrated strong performance on standard video prediction benchmarks and are widely regarded as effective in modeling temporal dependencies and spatial feature organization. However, unlike natural video sequences where temporal variation is dominated by local texture dynamics and intensity changes, swarming colony expansion is driven primarily by the propagation of morphological fronts, the generation and transmission of local geometric perturbations, and the temporal evolution of the global expansion velocity field. Consequently, the ability to accurately track front curvature, scale variation, and perturbation continuity is far more critical than minimizing pixel–wise reconstruction error. To ensure a fair comparison, all models were retrained from scratch under identical data partitions, observation/prediction splits, and evaluation protocols, and their outputs were evaluated using the mIoU, HD95, and ASSD metrics, as shown in Fig. 2.3a.
The results indicate that existing video prediction models exhibit clear limitations in this forecasting setting. MIM and SimVP+gSTA achieve relatively high region–level overlap, with mIoU of 89.32% and 90.52%, respectively, suggesting that they are able to preserve the overall expansion pattern to some extent. However, both models perform suboptimally in terms of HD95 and ASSD, indicating persistent discrepancies between the predicted and true colony fronts. In particular, PredRNN and PredRNNv2, although designed to enhance temporal memory and long–range dependency modeling, show consistently higher HD95 and ASSD values than the other models. This suggests that their hidden–state–accumulation–based recurrent formulation is susceptible to error amplification during extrapolation, making it difficult to maintain accurate front localization and stable representation of local geometric details over extended prediction horizons. In addition, SimVP+TAU demonstrates moderate performance across all three metrics, failing to strike an effective balance between region–level consistency and boundary precision.
In contrast, Morpher achieves the best performance across all evaluation metrics, with a mIoU of 95.42%, an HD95 of 10.61 pixels and an ASSD of 3.93 pixels, representing a 5.4% gain in region–level overlap and reductions of 42.0% and 55.7% in HD95 and ASSD relative to the strongest baseline, SimVP+gSTA. Although only the final 20% of each sequence requires extrapolative prediction, this window lies in the late stage of colony expansion, where the advancing interface is large and highly sensitive; even small deviations grow into disproportionately large discrepancies as the front progresses. The marked decreases in HD95 and ASSD show that Morpher not only aligns with the true expansion pattern but also preserves the geometry of the propagating front as it deforms. Each percentage point of mIoU thus corresponds to a tangible gain in propagation stability and in the maintenance of morphological integrity. By retaining the trajectory of the advancing edge with greater fidelity, Morpher produces forecasts that are more consistent with the underlying biological process of swarming expansion and capture features that matter for nutrient access, epithelial protection and microbial competition.
Figure 4: Failure modes of generic video prediction in swarming morphology forecasting and the role of morphology-aware representation. a, Performance comparison under an 80% observation / 20% prediction protocol. State-of-the-art video prediction models are evaluated under identical input–output settings for long-horizon forecasting of swarming colony expansion. While several methods attain reasonable region-level overlap, boundary-related errors remain pronounced across models. Morpher achieves higher mIoU together with consistently lower HD95 and ASSD, reflecting improved accuracy in front propagation and boundary geometry. b, Long-horizon qualitative forecasts on two representative sequences spanning distinct growth regimes. In the dense–finger phase (top), generic predictors either progressively smooth high-curvature fingers or fail to follow the true front dynamics, leading to increasing deviation from the ground truth. In the late-stage near–concentric ring expansion (bottom), maintaining global radial scale and subtle curvature becomes challenging for these models. Across both regimes, Morpher maintains a coherent and temporally consistent front geometry over extended prediction horizons, remaining closer to the true front. c, Effect of segmentation backbone on forecasting performance. Using TexPol–Net masks leads to more stable and accurate final-frame forecasts, whereas YOLOv12 [41], the strongest alternative backbone in this task, produces competitive results with residual boundary deviation and curvature attenuation. In the final predicted frame, TexPol–Net achieves higher IoU than YOLOv12 across both growth regimes, with improvements from 94.21% to 96.63% in the dense–finger phase and from 93.30% to 96.38% in the near–concentric ring expansion.
To further corroborate the quantitative results, we compare the predicted colony evolution across representative growth stages, as illustrated in Fig. 2.3b. During the dense–finger–instability–dominated expansion phase, MIM displays a clear tendency toward morphological smoothing over time: fine–scale protrusions at the advancing front gradually collapse into low–curvature boundaries, suppressing the inherent radial anisotropy and attenuating the transmission of local perturbations, which in turn compromises the accuracy of global expansion characterization over extended prediction horizons. The SimVP variants employing TAU or gSTA temporal units preserve the overall outline of the colony to a certain extent, but fail to maintain the dense–finger morphology, frequently exhibiting truncated lobe tips, blurred curvature transitions, and unstable front propagation rates; their predicted boundaries also become progressively rounded, reflecting insufficient geometric resolution and resulting in noticeable boundary drift. MAU and the recurrent architectures PredRNN and PredRNNv2 generate shapes that are superficially close to the ground truth in terms of global geometry, yet their temporal evolution is consistently slower, characterized by delayed dense–finger development, systematic underestimation of front propagation speed, and reduced amplitude of perturbation transmission, producing a temporal mismatch in both radial scale and finger–scale development. In contrast, Morpher yields stable and coherent predictions across growth regimes, maintaining both the spatial continuity of local perturbations and the global advancement and curvature modulation of the expanding front, while preserving a stable advancing interface without loss of curvature transmission or propagation continuity.
In the second sequence, the colony enters a late–stage expansion regime with a predominantly disc–like symmetry and a stabilized advancing front. In this context, the forecasting challenge shifts from preserving dense–finger morphology to maintaining the global radial scale and subtle boundary curvature modulation. MAU, MIM, PredRNN, and PredRNNv2 consistently underestimate the overall expansion radius, producing noticeably contracted predictions. This suggests that their temporal representations favor local smoothing and are unable to capture the long–range coordinated forces driving the sustained outward advancement of the expansion front and its slow, anisotropic curvature modulation. The SimVP model employing the gSTA temporal module better approximates the global scale, yet its predicted boundaries remain irregular and exhibit gradual drift over time, indicating limited geometric stability once high–frequency edge cues have diminished. The SimVP model using the TAU unit performs intermediately, but displays inconsistent boundary curvature and localized radial deviations across extended prediction intervals, reflecting difficulties in balancing temporal smoothness with geometric precision. In contrast, Morpher maintains both the disc–like radial organization and the fine–scale curvature variations throughout this stabilized growth phase. Its predicted boundaries closely follow the true expansion front without contraction, drift, or distortion, demonstrating that Morpher preserves stable global organization even when local structural cues weaken. Together with the observations from the dense–finger–dominated regime, these results indicate that Morpher captures the underlying evolutionary dynamics of swarming colonies rather than relying on short–term motion extrapolation or local texture continuity.
The limited performance of conventional video prediction models in this task stems from their emphasis on global pixel–level variation rather than on the dynamical organization of the expanding morphological front. Morpher departs from this paradigm by operating directly on the binary mask sequences produced by TexPol–Net, enabling temporal modeling to take place within a compact latent representation that encodes structural configuration rather than raw appearance. Within this latent space, Morphon works together with the autoregressive feedback loop to continuously update the representation based on the accumulated history of colony growth. This mechanism allows Morpher to directly capture the speed, directional bias, and spatial coordination of front propagation. Coupled with a multi–scale decoder that preserves continuity between successive morphological states, the model maintains coherent global organization and smooth boundary evolution even over long forecast horizons. As a result, Morpher produces long–range predictions that exhibit higher dynamical consistency and more faithful morphological progression than existing video prediction approaches.
2.4 Small improvements in front recovery yield large gains in forecasting
While the predictive performance of Morpher demonstrates its ability to capture temporal dependencies, the reliability of long–term forecasts hinges on the accuracy of the input representation—the segmentation masks delineating colony morphology. To assess this relationship, we compared Morpher coupled with our TexPol–Net segmentation backbone to the same forecasting framework driven by YOLOv12. YOLOv12 achieves the second–highest segmentation accuracy on our swarming colony dataset—second only to TexPol–Net—and thus serves as the strongest alternative baseline for this task. This comparison provides a clear and controlled evaluation of how the segmentation backbone affects subsequent temporal modeling. Both systems follow the identical two–stage pipeline illustrated in Fig. 1, using the same Morpher architecture, training settings, and evaluation criteria; the only difference lies in the source of the input segmentation masks.
Although YOLOv12 (91.81 mAP50:95) trails TexPol–Net (92.48 mAP50:95) by less than one percentage point in static segmentation accuracy, this seemingly minor gap translates into a disproportionately large effect on temporal forecasting. As illustrated in Fig. 2.3c, the slight boundary inaccuracies present in YOLOv12 masks accumulate through Morpher’s autoregressive updates, leading to a gradual drift of the colony front and a measurable loss of curvature fidelity over time—up to a 2.4–3.1 IoU-point gap after 20% forecasting. Forecasts initialized from YOLOv12 masks exhibit smoothed protrusions and reduced dense–finger morphology, consistent with progressive attenuation of high–frequency boundary features. In contrast, TexPol–Net masks preserve sharper front geometry and coherent internal texture cues, allowing Morpher to maintain stable propagation of the expanding rim with minimal geometric distortion.
Temporal coherence in morphological forecasting depends not only on the expressiveness of the temporal model but also on the structural fidelity of its spatial inputs. In swarming dynamics—where front curvature, anisotropy, and the organization of dense fingers jointly define the expansion process—even slight degradations in segmentation accuracy can disrupt the causal progression linking successive morphological states. TexPol–Net mitigates this sensitivity through its TEA and PCA mechanisms, which together sharpen fine–scale front delineation while preserving the global radial organization of the colony. These inductive priors yield geometrically consistent masks that capture the underlying biophysical architecture with high fidelity, enabling Morpher to propagate colony morphology smoothly and without structural drift. From a modeling standpoint, the results reveal that even a seemingly marginal one–percentage–point improvement in segmentation accuracy can induce a qualitative shift in system behavior—from cumulative error amplification to dynamical stability. The integration of TexPol–Net and Morpher thus forms a synergistic framework that unites spatial precision with temporal reasoning, enabling quantitative reconstruction of curvature–driven, long–range evolution in swarming fronts and, more broadly, extending predictive modeling to other forms of collective biological motion.
2.5 Temporal models and inference strategies for front forecasting
We used Morpher to systematically probe how temporal sequence models, inference strategies, and the Morphon module influence long–horizon forecasts of swarming colony morphology, as summarized in Fig. 2.5b. Four architectures—a simple RNN, GRU, LSTM, and Transformer encoder—were evaluated under parallel versus autoregressive decoding, with or without Morphon.
Forecasting quality was assessed using a set of complementary metrics capturing geometric accuracy and dynamical stability, including region–level overlap, boundary deviation, front–propagation dynamics, temporal consistency, and angular organization. The quantitative results underlying these comparisons are reported in detail in Table 2, which provides the full ablation statistics for each architectural and inference configuration.
A first observation is that autoregressive decoding consistently produced more faithful colony evolution than parallel one–shot decoding. Across architectures, switching from parallel to autoregressive inference improved either mIoU, boundary metrics, or both. For example, the Transformer with Morphon achieved 94.80% mIoU and an ASSD of 4.63 px in parallel mode, compared with 95.42% mIoU and 3.93 px ASSD under autoregressive decoding. Similar patterns were observed for GRU and LSTM, where autoregression reduced HD95 and ASSD and kept RMSE at low values. This behavior matches the biology of the system: swarming expansion proceeds through many small, incremental deformations of the front, and each future shape is tightly constrained by the immediately preceding configuration. Parallel decoding enforces all future frames in a single pass, which tends to collapse the diversity of plausible growth trajectories into a smoother, more symmetric contour and suppresses local protrusions. Conditioning each prediction on the previous frame allows the model to adjust the advancing front step by step, preserving fingering, local curvature, and realistic front propagation over long horizons.
@@ 2.5 Temporal models and inference strategies for front forecasting
Figure 5: Morpher enables geometry-consistent long-horizon forecasting of swarming colony morphology. a, Architecture of Morpher. Observed segmentation masks are mapped to a compact morphological latent sequence by a multi-scale spatial encoder. Future evolution is predicted autoregressively in latent space, with each decoded morphology fed back to condition subsequent steps. The Morphon module selectively retrieves informative past states via cross-attention and integrates them through a learnable gate, while a multi-scale decoder reconstructs spatial detail and preserves peripheral protrusions and front curvature. b, Systematic evaluation of temporal architectures and inference strategies under a fixed 80% observation / 20% prediction protocol. Across eight complementary metrics capturing geometric accuracy and dynamical stability, autoregressive inference tends to outperform parallel decoding, and the inclusion of Morphon improves boundary fidelity and temporal consistency across sequence models (RNN, LSTM, GRU and Transformer). c, Robustness of morphological forecasting to observation length. For autoregressive models with Morphon enabled, prediction accuracy and stability improve smoothly as the observation ratio increases from 50% to 90%, indicating that Morpher exploits additional temporal context without overfitting to a specific observation–prediction split.
The sequence models themselves formed a clear performance hierarchy that aligns with their representational capacity. The plain RNN was consistently weakest on geometry-sensitive metrics. Without Morphon in parallel mode, it reached 93.23% mIoU but suffered from large boundary errors (ASSD 6.02 px, $|\Delta\mathrm{NAS}|$ 19.54%), yielding overly round and contracted colonies. Even with autoregression and Morphon, ASSD remained above 4.3 px and anisotropy above 16%, indicating that local perturbations at the front were rapidly forgotten and the contour relaxed toward a smooth disk. Introducing gating improved this behavior. The GRU in parallel mode without Morphon already reduced ASSD to 5.17 px and $|\Delta\mathrm{NAS}|$ to 15.31%, and in its best configuration (autoregressive with Morphon) it further lowered ASSD to 4.06 px and achieved the smallest RMSE (2.06 px per frame) among all models, indicating particularly sharp tracking of front advancement and rapid local changes. The LSTM occupied an intermediate regime: it produced lower HD and HD95 than the RNN and stable RMSE values (2.26–2.84 px per frame), capturing global expansion reliably but reacting less strongly to fine, transient undulations of the front.
The Transformer–based Morpher provided the most balanced overall forecasts. In the autoregressive Morphon setting, it achieved the highest mIoU (95.42%), the lowest ASSD (3.93 px), and the smallest anisotropy deviation ($|\Delta\mathrm{NAS}|=13.13\%$), while keeping both the RMSE of front propagation and $|\Delta\mathrm{H}_{2}|$ at low values. These scores indicate that the model not only captured the colony’s radial expansion and average outline, but also maintained fingering, sectoring, and higher–order harmonic structure. Self–attention over the encoded history enables the Transformer to couple distant parts of the front and integrate long–range constraints, which is particularly important for reproducing radial symmetry–breaking and coordinated fingering across the colony.
Morphon provided an additional axis of improvement by encoding a persistent structural memory of past fronts. For every architecture and inference strategy, enabling Morphon increased mIoU and sharpened the boundary. For instance, in parallel mode the RNN improved from 93.23% to 94.22% mIoU and reduced ASSD from 6.02 px to 4.85 px when Morphon was switched on. Autoregressive GRU improved HD95 from 12.03 px without Morphon to 10.14 px with Morphon, and the Transformer under autoregression decreased ASSD from 5.26 px to 3.93 px when Morphon was enabled. Angular metrics showed the same tendency: in the strong Transformer configuration, $|\Delta\mathrm{NAS}|$ decreased from 16.83% without Morphon in parallel mode to 13.13% with Morphon in autoregressive mode, reflecting the combined effect of autoregression and Morphon in reinforcing angular fidelity. Across models, $|\Delta\mathrm{H}_{2}|$ remained below 2.0%, indicating that second–harmonic structure was generally well preserved, while Morphon primarily helped to maintain the finer anisotropic patterning of the front.
Accurate forecasting depends on both an appropriate inference paradigm and biologically informed memory. Autoregressive decoding is crucial for maintaining causal continuity of growth, while Morphon stabilizes curvature and anisotropy by carrying forward structural cues. Within this design, the Transformer–based Morpher offers the best compromise between boundary fidelity, propagation dynamics, and angular organization. The GRU emphasizes rapid front dynamics, the LSTM ensures smooth global expansion at the cost of some fine structure, and the plain RNN underestimates both fingering and symmetry breaking. These patterns support a mechanistic interpretation: models that can remember and attend to past colony structure are best suited to predict how microbial collectives remodel themselves in time.
2.6 Forecasting stability across observation ratios
We next examined how Morpher behaves as the amount of observed history is systematically varied. For each temporal backbone operating in its best–performing configuration (autoregressive decoding with Morphon), the observation fraction was increased from 50% to 90%. Model performance was then evaluated using the same eight metrics described above. The resulting trends are summarized in Fig. 2.5c, with the corresponding quantitative values reported in Table 3. This analysis tests whether Morpher is tuned to a narrow observation–prediction split or instead maintains stable forecasting behavior across a broad range of experimental protocols.
Geometry-based metrics showed a smooth and largely monotonic improvement as more of the past was revealed. For all four sequence models, mIoU increased from the high–80% range at 50% observation to above 96% at 90%, while HD95 and ASSD decreased correspondingly. The Transformer, for example, improved from 88.22% to 96.79% mIoU and reduced ASSD from 9.49 px to 2.75 px as the observation ratio increased from 50% to 90%. The relative ranking of sequence models remained essentially unchanged: at each observation level the Transformer offered the best compromise between overlap and boundary error, GRU and LSTM occupied an intermediate regime, and the plain RNN lagged slightly behind. This preservation of ordering indicates that the choice of temporal architecture is more important than the precise observation ratio for recovering the colony outline and front position.
Front–propagation dynamics, measured by RMSE of radial expansion rate, exhibited a similar saturation pattern. The largest gains occurred between 50% and 60% observation, after which RMSE values fluctuated within a narrow band of roughly 2.0–2.3 px per frame across models. In other words, once the model has seen slightly more than half of the trajectory, the incremental benefit of additional frames for predicting front–propagation speed is modest. TCI behaved in a comparably stable way but with slightly larger relative variation. TCI values clustered around 63–65% from 50% to 80% observation, with the LSTM reaching the highest TCI (65.32%) at 80%. At 90% observation the prediction window becomes too short for a meaningful fluctuation index, which is why TCI is not reported there. From a mathematical standpoint, both RMSE and TCI are based on temporal differences rather than static shapes; they are sensitive to small phase shifts in the timing of front accelerations and pauses, even when the underlying boundary is well aligned.
Angular metrics, particularly $|\Delta\mathrm{NAS}|$, varied more across observation ratios and did not follow a simple monotonic trend. For the Transformer, $|\Delta\mathrm{NAS}|$ decreased from 14.62% at 50% observation to 11.93% at 60%, then increased to 15.57% at 70% before dropping again to 13.13% at 80% and settling around 14% at 90%. The recurrent models showed the same qualitative behaviour, with a pronounced peak near 70–80%. This seemingly counterintuitive pattern—more observed history not always yielding smaller anisotropy error—arises from how NAS is defined: it aggregates angular contrasts over the entire prediction horizon. Increasing the observation ratio shortens this horizon and shifts it toward later times, where the colony has already developed complex sectoring and branch splitting. In this regime, even small local disagreements in when a branch emerges or merges can produce relatively large changes in the angular power spectrum. Mathematically, $|\Delta\mathrm{NAS}|$ behaves more like a measure of angular fluctuation intensity than a simple bias; it is highly sensitive to local phase errors of individual fingers, and therefore displays a non–monotonic dependence on the observation window.
The second–harmonic descriptor $|\Delta\mathrm{H}_{2}|$ showed a different but equally non–trivial trend. For the three recurrent architectures, $|\Delta\mathrm{H}_{2}|$ increased almost monotonically from $\sim\!1.1$–1.2% at 50% observation to around 2.2–2.3% at 90%. The Transformer followed the same overall tendency, with values around 1.18–1.35% at 50–60% observation and $\sim\!2.2\%$ at 90%, albeit with a local maximum at 70% and a mild dip at 80%. At first glance, this appears to contradict the intuition that more context should always improve harmonic structure. However, $|\Delta\mathrm{H}_{2}|$ isolates only the second angular mode of the front. As the colony matures, curvature energy is progressively redistributed into higher–order modes through branch splitting and tip instabilities. The prediction window at high observation ratios is therefore dominated by subtle rearrangements of fine–scale protrusions rather than by the initial quadrupolar shape. In this late regime, tiny mismatches in the amplitude or phase of low–frequency modes can produce larger absolute differences in the second harmonic, even when the full boundary and overall anisotropy are well captured. From a mathematical perspective, the metric is probing how accurately the model tracks a single Fourier component in a spectrum whose energy is shifting across modes; as the spectrum becomes more crowded, that task becomes harder rather than easier.
Morpher remains stable across observation ratios for coarse geometric descriptors and front–propagation speed, whereas the more sensitive temporal and angular metrics reveal where the remaining difficulty lies. Once roughly 60–70% of the trajectory has been observed, mIoU, HD–based distances, ASSD, and RMSE change only gradually with further increases in observation, and the relative ranking of sequence models remains fixed. In contrast, $|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$ respond strongly to how the forecasting window slices the late–time dynamics of sectoring and finger rearrangement, whereas TCI shows modest but noticeable sensitivity. This pattern indicates that changing the observation–prediction split does not alter the relative ranking of the sequence models in recovering global expansion, and that the main residual challenges are concentrated in the most fluctuation– and angle–sensitive aspects of the swarming front.

### v2

@@ 2.3 Segmentation defines the predictive state
2.4 Limits of appearance-based prediction
Morphology-level forecasting is ultimately judged by whether the active front is localized and whether its fine branches are preserved. Region overlap (mIoU) can remain high while the contour drifts; we therefore treat boundary-sensitive distances (HD95 and ASSD) as the primary indicators of predictive fidelity, and use overlap as a complementary check on coarse extent (Figure 4a).
We compared Morpher with representative video prediction models, including MAU [9], MIM [63], PredRNN [61], PredRNNv2 [62], the original SimVP implementation with the TAU temporal unit [18, 57], and the improved SimVPv2 variant with the gSTA module [56], all retrained under identical splits and evaluated on the same 80% observation / 20% prediction protocol (Figure 4a). This ordering holds across all metrics. MIM and SimVP+gSTA preserve coarse regional extent to some degree, reaching mIoU values of 89.32% and 90.52%, but remain substantially worse in boundary accuracy. PredRNN and PredRNNv2 exhibit still larger HD95 and ASSD, indicating stronger drift during extrapolation. SimVP+TAU remains intermediate without resolving the overlap–boundary tradeoff.
Figure 4: Failure modes of generic video prediction in swarming morphology forecasting and the role of morphology-aware representation. a Performance under an 80% observation / 20% prediction protocol. Morpher achieves the best overall accuracy (95.42% mIoU, 10.61 px HD95, 3.93 px ASSD), indicating improved front localization and boundary fidelity. b Long-horizon forecasts on two representative sequences. Generic predictors smooth finger-like branches or accumulate boundary drift, whereas Morpher maintains a coherent front closer to the ground truth across regimes. c Effect of segmentation backbone on forecasting. TexPol–Net masks yield more accurate final-frame predictions than YOLOv12 [58], increasing IoU from 94.21% to 96.63% in the anisotropic branching regime and from 93.30% to 96.38% in the near-concentric regime.
Morpher performs best on all three metrics, reaching 95.42% mIoU, 10.61 px HD95, and 3.93 px ASSD. Relative to the strongest baseline, SimVP+gSTA, this corresponds to a 5.4% gain in overlap and reductions of 42.0% and 55.7% in HD95 and ASSD. These gains are obtained under an 80% observation / 20% prediction protocol, where prediction is restricted to the late expansion stage, when the front is already extended and highly sensitive to small geometric deviations.
The qualitative failures follow the same ordering (Figure 4b). In the anisotropic branching regime, MIM progressively smooths protrusions; SimVP variants retain a coarse outline but truncate lobe tips and blunt high-curvature sectors; MAU, PredRNN, and PredRNNv2 lag in propagation, yielding fronts that are spatially plausible but temporally delayed. In the late near-concentric regime, the same models underestimate radial extent or accumulate boundary drift. Morpher remains closest to the true front across both regimes, preserving local perturbations without losing global scale. The separation reflects a mismatch in what is being modeled: generic video predictors favor appearance continuity, whereas Morpher treats morphology itself as the evolving state.
2.5 Segmentation errors accumulate during forecasting
A causal link between measurement fidelity and forecast stability can be tested by a controlled swap: holding the temporal model fixed while changing only the segmentation backbone that defines the state. We therefore paired the same Morpher architecture with either TexPol–Net or YOLOv12. The static difference is small: TexPol–Net achieves 92.48% mAP50:95, whereas YOLOv12 reaches 91.81%. The temporal consequence is not. Over the final 20% prediction window, YOLOv12 masks incur a 2.4–3.1 IoU-point loss in final-frame prediction.
This amplification is visible in both representative regimes (Figure 4c). In the anisotropic branching sequence, final-frame IoU increases from 94.21% to 96.63%; in the near-concentric sequence, from 93.30% to 96.38%. This establishes a causal amplification mechanism: sub-percent-level segmentation differences are sufficient to induce multi-point degradation in long-horizon prediction. The mechanism is intrinsic to autoregressive boundary evolution, in which small geometric biases are recursively fed back as state and thereby converted into cumulative drift.
2.6 Temporal modeling of front dynamics
Temporal prediction must preserve two coupled properties: where the boundary lands (geometric fidelity) and how it gets there (dynamical consistency). The evaluation therefore spans eight complementary metrics (Figure 5b), with a deliberate hierarchy for interpretation: boundary-sensitive distances (HD, HD95, ASSD) and overlap (mIoU) report localization; propagation RMSE and TCI test trajectory-level stability; angular deviations ($|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$) diagnose whether anisotropic branching is preserved beyond coarse extent.
@@ 2.6 Temporal modeling of front dynamics
Figure 5: Morpher enables geometry-consistent long-horizon forecasting of swarming colony morphology. a Morpher architecture. Observed masks are encoded into a compact morphological latent sequence, and future evolution is predicted autoregressively with decoded states recursively conditioning subsequent steps. The Morphon module retrieves past states via cross-attention and integrates them through a learnable gate, while a multi-scale decoder preserves boundary detail and front geometry. b Evaluation of temporal architectures and inference strategies under an 80% observation / 20% prediction protocol. Autoregressive inference consistently outperforms parallel decoding, and Morphon improves boundary fidelity and temporal consistency across RNN, LSTM, GRU, and Transformer models. c Robustness to observation length. With Morphon enabled, prediction accuracy and stability improve as the observation ratio increases from 50% to 90%, indicating effective use of temporal context without overfitting.
We analyzed temporal architectures (RNN, GRU, LSTM, and Transformer) under parallel and autoregressive decoding, with or without Morphon (Figure 5b). Autoregressive decoding consistently outperforms parallel prediction. For Transformer with Morphon, performance improves from 94.80% mIoU and 4.63 px ASSD in parallel mode to 95.42% and 3.93 px under autoregressive decoding. GRU and LSTM show the same directional improvement. Stepwise state updates preserve local geometric continuity, whereas one-shot prediction smooths the front.
The backbones nonetheless separate along different axes of performance. The plain RNN is weakest: without Morphon in parallel mode, it reaches 93.23% mIoU, 6.02 px ASSD, and $|\Delta\mathrm{NAS}|=19.54\%$, producing overly round and contracted fronts. GRU is substantially stronger; even without Morphon in parallel mode, it reduces ASSD to 5.17 px and $|\Delta\mathrm{NAS}|$ to 15.31%, and in its best configuration it further lowers ASSD to 4.06 px and achieves the lowest propagation RMSE, 2.06 px per frame. LSTM is more stable than RNN and keeps RMSE between 2.26 and 2.84 px per frame, but reacts less strongly than GRU to fine front undulations.
Critically, the model that minimizes propagation error is not the one that best preserves morphology. GRU achieves the lowest RMSE (2.06 px per frame). Under the same setting, it remains worse than Transformer in boundary fidelity and anisotropic structure preservation. In the autoregressive Morphon setting, Transformer yields the highest mIoU (95.42%), the lowest ASSD (3.93 px), and the smallest anisotropy deviation ($|\Delta\mathrm{NAS}|=13.13\%$), while keeping both propagation RMSE and $|\Delta\mathrm{H}_{2}|$ low. Accurate front advance and faithful morphology reconstruction are therefore not identical objectives. The former is largely a gated-memory problem, whereas the latter requires long-range retrieval of branch history and spatial configuration.
Morphon improves every backbone, and its effect is diagnostic rather than cosmetic: it specifically targets the non-Markovian part of boundary evolution, where similar instantaneous curvatures can lead to different futures depending on earlier branch history. In parallel RNN, Morphon raises mIoU from 93.23% to 94.22% and reduces ASSD from 6.02 to 4.85 px. In autoregressive GRU, it reduces HD95 from 12.03 to 10.14 px. In Transformer, it lowers ASSD from 5.26 to 3.93 px and decreases $|\Delta\mathrm{NAS}|$ from 16.83% to 13.13%. Notably, $|\Delta\mathrm{H}_{2}|$ remains below 2.0% across models, suggesting that low-order harmonic structure is largely preserved even by weaker predictors. The main separation therefore lies not in coarse global shape, but in the preservation of finer anisotropic organization along the advancing front.
2.7 Scaling with observation length
The observation ratio was varied from 50% to 90% for each backbone in its best configuration, namely autoregressive decoding with Morphon (Figure 5c). We begin at 50% observation because the early phase is morphologically convergent across colonies (Figure 6d–f), and prediction only becomes well-posed once colony-specific geometry begins to diverge. Geometry-based metrics improve smoothly with longer history. For the Transformer, mIoU rises from 88.22% at 50% observation to 96.79% at 90%, while ASSD falls from 9.49 to 2.75 px. The relative ordering of the backbones remains unchanged across the full range, indicating that temporal architecture matters more than the exact observation–prediction split.
Front-propagation RMSE follows the same direction but saturates earlier. The largest gain appears between 50% and 60% observation; beyond that, values remain within a narrow 2.0–2.3 px per frame band. TCI is similarly stable, clustering around 63–65% from 50% to 80% observation, with the highest value observed for LSTM at 80% observation (65.32%). At 90%, the prediction window is too short for a stable TCI estimate.
Angular metrics expose a stricter regime and reveal a second constraint: more history does not guarantee better anisotropy preservation because the error is dominated by phase alignment of late-stage branch rearrangements. For the Transformer, $|\Delta\mathrm{NAS}|$ changes from 14.62% at 50% observation to 11.93% at 60%, rises to 15.57% at 70%, drops to 13.13% at 80%, and remains around 14% at 90%. $|\Delta\mathrm{H}_{2}|$ shows a different but equally late-stage-sensitive trend: for the Transformer it is 1.18–1.35% at 50–60% observation and approaches $\sim$2.2% at 90%. Additional history therefore yields diminishing returns for low-order shape statistics, while leaving phase-sensitive alignment of late front rearrangements as the dominant residual difficulty.
2.8 Generalization of trajectory-level dynamics
Out-of-sample stability was tested by leave-one-out cross-validation, evaluating each trajectory with models trained on all remaining trajectories (Figure 6). Performance approaches saturation as the training set expands, with only marginal gains once most trajectories are included (Figure 6a), indicating that the dominant modes of swarming variation relevant for forecasting are already well represented.
Figure 6: Generalization, accuracy, and dynamical consistency of swarming morphology forecasting under data-limited conditions. Results are evaluated using leave-one-out cross-validation across 81 independent colonies. a Effect of training set size on forecasting performance (80% observation / 20% prediction). Metrics converge with increasing data, indicating that the observed variability is sufficiently captured. b Static forecasting accuracy (80% / 20%) measured by region- and boundary-based metrics. c Dynamical consistency (80% / 20%) relating propagation error, anisotropy deviation, and normalized performance. d–f Time-aligned trajectories of effective radius, perimeter, and front propagation velocity under a 50% observation / 50% prediction protocol, with individual trajectories and mean trends. g Prediction-horizon–aligned trajectories (50% / 50%) showing forecast evolution within the prediction window. h Prediction-horizon–aligned trajectories (80% / 20%) showing consistent behavior across observation regimes.
Leave-one-out evaluation separates static accuracy from dynamical fidelity. Region- and boundary-based metrics remain concentrated in a high-performance regime (Figure 6b), whereas dynamical measures separate propagation error, anisotropy deviation, and harmonic distortion more clearly (Figure 6c). The model therefore generalizes well in mask overlap without collapsing the more sensitive dynamical dimensions into the same signal.
This separation is preserved at the trajectory level. Under a 50% observation / 50% prediction protocol, predicted effective radius and perimeter closely track the true time-aligned trajectories, and front propagation velocity reproduces the transition from rapid early expansion to later stabilization (Figure 6d–f). Agreement remains after alignment by prediction horizon (Figure 6g). Under an 80% observation / 20% prediction split, the same dynamical organization is retained (Figure 6h): longer observation tightens prediction over a shorter horizon without changing the overall structure of the trajectories. Across settings, Morpher preserves not only colony extent but also the temporal organization of expansion.